OpenStudy (anonymous):

The number of vertical tangents to the graph of y^2=x-x^3 is? Please guide me through.

OpenStudy (solomonzelman):

The number of vertical tangent to \(\large\color{#000000 }{ \displaystyle y^2=x-x^3 }\) will be the same as the number of vertical tangents to \(\large\color{#000000 }{ \displaystyle y=\sqrt{x-x^3} }\) Differentiate the function, and the points where the derivative is undefined are going to be vertical tangents.

OpenStudy (solomonzelman):

find the derivative, and set the denominator of the derivative =0.